OpenStudy (anonymous):

How much force is needed to stop a 100-kg football player if he decelerates at 20 m/s²? A. 20 N B. 2000 N C. 200 N D. 5 N

OpenStudy (anonymous):

@timaashorty

OpenStudy (anonymous):

@aaronq

OpenStudy (whpalmer4):

\[F=ma\] You know \(m\), you know \(a\) \[1 ~kg\cdot\frac{m}{s^2} = 1~N\]

OpenStudy (whpalmer4):

By the way, this really would be better posted in the Physics section...